Suppose that an ion source in a mass spectrometer
Answer #1

given Kinetic energy k = dV*q


0.5*m*v^2 = dV*q

v = sqrt(2*dV*q/m) = sqrt((2*2.2*10^3*2*1.6*10^-19)/(3.27*10^-25))


v = 6.56*10^4 m/s

In magnetic field


Fb = Fc


q*v*B = m*v^2/r

r = mv/qB


r = (3.27*10^-25*6.56*10^4)/(2*1.6*10^-19*0.71)

r = 0.0944 m     <<<<<,------answer